After researching the competitors of EJH​ Enterprises, you determine that most comparable firms have the following valuation​ ratios:

Comp 1

Comp 2

Comp 3

Comp 4

​EV/EBITDA

1212

1111

12.512.5

1010

​P/E

1919

1818

2020

1717

EJH Enterprises has EPS of

​$1.801.80​,

EBITDA of

​$295295

​million,

​$2525

million in​ cash,

​$4444

million in​ debt, and

103103

million shares outstanding. What range of prices is consistent with both sets of​ multiples?

The range of prices will​ be:

Lowest price within both​ ranges, the​ P/E and ​ EV/EBITDA ranges, is

​$nothing.

  ​(Round to two decimal​ places.)Highest price within both​ ranges, the​ P/E and the​ EV/EBITDA ranges, is

​$nothing.

​(Round to two decimal​ places.)
Answer #1

We need to calculate the lowest price using both multiples.

Using lowest P/E, Price = P/E x EPS = 17 x 1.8 = 30.60

Using lowest EV/EBITDA, EV = 10 x 295 = 2950

Equity Value = EV - Debt + Cash = 2950 - 44 + 25 = 2,931

Lowest Price = 2,913 / 103 = 28.46... which is the lowest price

Similarly, using highest P/E, Price = 20 x 1.8 = 36

Highest EV, EV = 12.5 x 295 = 3,687.5

Highest Price = (2,687.5 - 44 + 25) / 103 = 35.62

Hence, $36.0 is the highest price.